# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

### v2

@@ 7 Discussion
We present a new framework to analyze the three main components relevant to weight imprinting, namely, weight generation, normalization, and aggregation. Within this framework, state-of-the-art imprinting strategies become special cases. This allows for a comprehensive analysis of different approaches through systematic experiments and leads us to generalize to a new, best-performing imprinting strategy. That is, using k-means weight generation with L2 normalizations and max aggregation outperforms all previously studied methods (see fig. 2).
@@
In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small. Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed in [4], highlighting its advantage in few-shot scenarios.
@@
The max aggregation directly scales with the magnitude of the weights. Normalization (NORMpost) ensures that all class weights contribute equally to the output. Nearest neighbor (l-nn) aggregation is not as affected by the lack of normalization, since it uses Euclidean distance. Although still part of common procedure, normalizations for embeddings (NORMpre and NORMinf) appear to have minimal impact on performance.
@@
While nearest neighbor (m-nn) aggregation (AGG) performs well when all data is saved (e.g., when there are no storage constraints), max aggregation with limited number of representative proxies (e.g., $k$-means) is an efficient alternative without a substantial loss in performance.
@@
During training, the last-layer weights of a FM collapse to their respective class means. This proves the success of mean imprinting on known classes. New, out-of-distribution data, however, often shows less collapse, making it beneficial to imprint more than one proxy.
@@
Since our experiments are limited to a small selection of foundation models and tasks, running additional experiments could strengthen the statistical significance. While imprinting alone provides an efficient solution to transfer learning, we do not investigate the benefit of combining it with optimization methods like gradient-based learning when more samples become available. This combination could use imprinting as initialization or apply metric learning to improve imprinting capabilities.
@@
The usage of both weight and activation sparsity as in [43] could change the within- and between-class variability in favor of using a higher number of proxies. Synaptic intelligence approaches like the weight saturation presented in this work are paths of further study. We use the penultimate layer embeddings for generating the classifier weights. An interesting area of study could be extracting embeddings from previous layers of the FM for this purpose. Lately, the study in [44] showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable. Apart from that, it would be interesting to imprint the weights of other layers as well (see, for example, [10]).
8 Conclusion
We investigated imprinting as an efficient method for transfer learning based on foundation models. Within our new framework, we found a new imprinting strategy that outperforms all previously studied ones. The phenomenon of neural collapse provides theoretical proof for its success.
Acknowledgments and Disclosure of Funding
Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528483508 - FIP 12, as well as the European Union under the grant project 101079894 (COMFORT - Improving Urologic Cancer Care with Artificial Intelligence Solutions). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. Furthermore, we would like to thank Viet Anh Khoa Tran for initial discussions about the neural collapse phenomenon.
@@ Appendix A Appendix
A.1 Additional Tables and Critical Difference Diagrams
We provide additional tables and critical difference (CD) diagrams that are referenced and put into context in the main paper.
GEN $k$ Avg. acc. %
k-means 5 89.61
mean 1 87.76
k-medoids 5 85.59
k-cov-max 5 82.82
k-random 5 76.40
k-fps 5 63.74
@@ A.1 Additional Tables and Critical Difference Diagrams
Figure 13: Benchmarking GEN mechanism for $k\leq 5$ across FMs and Ts. Best NORM combination for each row is used implicitly. AGG is fixed to max. CD diagram depicts statistical significance of k-means as GEN. See fig. 5 for $k\leq 20$.
NORMinf NORMpre NORMpost Avg. acc. %
L2 none L2 87.76
none none L2 87.76
L2 L2 L2 87.72
L2 L2 quantile 80.70
L2 none quantile 80.71
none none quantile 80.71
L2 L2 none 84.04
L2 none none 73.75
none none none 73.75
Table 2: Benchmarking NORM across FMs and Ts shows crucial effect of L2 normalization. GEN is fixed to mean and AGG to max. See fig. 14 for average ranks.
@@
Figure 14: CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as "NORMinf & NORMpre & NORMpost". See table 2 for further details and average accuracies.
NORMinf NORMpre NORMpost Avg. acc. %
L2 none L2 91.44
none none L2 91.44
L2 L2 L2 91.48
L2 L2 quantile 90.86
L2 L2 none 89.99
L2 none quantile 80.36
none none quantile 80.36
L2 none none 79.22
none none none 79.22
Table 3: Benchmarking NORM across FMs and Ts. GEN is fixed to k-means with $k=20$ and AGG to max. See fig. 15 for average ranks.
@@
Figure 15: CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as "NORMinf & NORMpre & NORMpost". See table 3 for further details and average accuracies.
@@
Figure 16: k-means with different values for $k$ in n-shot scenarios with focus on $10\leq n\leq 400$. $95\%$ confidence intervals shown in shaded colors. Other variables are fixed to our previously described best method of using L2 for NORM and max as AGG. Note that only data for the meaningful case of $k\leq n$ is shown. It can be inferred that for MNIST and FashionMNIST, mean is not the best strategy anymore starting at roughly 50 samples. See fig. 10 for more values of $n$.
A.2 Differences between Foundation Models
@@
While an in-depth comparison of foundation models is beyond the scope of this paper, we believe it is important to highlight key observations. In particular, fig. 11 shows significantly lower $\mathcal{NC}_{1}$ scores for vit_b_16 and swin_b on their pre-training ImageNet data compared to the resnet models. We hypothesize that this difference is primarily due to model size and training regimes. The Transformer-based architectures (vit_b_16 and swin_b) have a considerably higher parameter count ($\approx 87$M) than the resnet models (11.7M and 25.6M, respectively). Additionally, vit_b_16 and swin_b were trained for three times as many epochs (300 vs. 90) while using a substantially lower learning rate (0.003 and 0.01 vs. 0.1). Notably, the embedding dimensions of these models are comparable, meaning that the observed differences in $\mathcal{NC}_{1}$ scores cannot be attributed to differences in representation dimensionality. Instead, we argue that the combination of larger model size, extended training duration, and lower learning rates likely contributes to greater overfitting, leading to more pronounced collapse.
@@
Figure 17: Averaged accuracy of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$ over number of proxies $k$ used for k-means in GEN. $95\%$ confidence intervals are shown in shaded colors. We set NORMpost and NORMinf to L2, and NORMpre to none. Accuracies of the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are shown in dotted lines. Besides the prominent peaks in accuracy at $k=d$ (as already observed in fig. 12), a consistent dip between $k=1$ and $k=d$ appears in Transformer-based models.
@@
Figure 17, similar to fig. 12, illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs). The key difference in this figure is the use of none for NORMpre instead of L2. This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig. 12, where L2 is used as NORMpre, and does not occur at all in the resnet models. We hypothesize that this difference arises from the distinct embedding distributions of CNN and Transformer architectures (see, e.g., [14, Figure S2]).

### v3

@@ 6 Conclusion
We present a new framework, IMPRINT, to analyze three main components relevant to weight imprinting, namely, weight generation, normalization, and aggregation. Within IMPRINT, state-of-the-art imprinting strategies become special cases. This allows for a comprehensive analysis of different approaches through systematic experiments and leads us to generalize to a new, best-performing imprinting variant. That is, using k-means weight generation with L2 normalizations and max aggregation, which outperforms all previously studied methods (see fig.˜2).
@@
In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small. Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed by qi-imprinting, highlighting its advantage in low-data regimes.
@@
The max aggregation directly scales with the magnitude of the weights. Normalization (NORMpost) ensures that all proxies have the same magnitude, preventing differences in vector norms from disproportionately affecting classifier predictions. Nearest neighbor (l-nn) aggregation is not as affected by the lack of normalization, since it uses Euclidean distance. Although still part of common procedure, normalizations for embeddings (NORMpre and NORMinf) appear to have minimal impact on performance.
@@
While nearest neighbor (m-nn) aggregation (AGG) performs well when all data is saved (e.g., when there are no storage constraints), max aggregation with limited number of representative proxies (e.g., through $k$-means) is an efficient alternative without a substantial loss in performance.
@@
During training, the last-layer weights of a FM tend to collapse to their respective class means. This proves the success of mean imprinting on known classes. New, out-of-distribution data, however, often shows less collapse, making it beneficial to imprint more than one proxy. Our results show that the advantage of using multiple proxies is strongly coupled to the degree of neural collapse: the accuracy gain from using $k>1$ proxies increases approximately log-linearly with $\mathcal{NC}_{1}$. In particular, multi-proxy imprinting yields substantial and predictable improvements on datasets with $\mathcal{NC}_{1}$ exceeding 1. While from a practical perspective, choosing the exact number of proxies $k$ based on pure greedy search with validation data (or as part of any AutoML pipeline) is still a valid option, our analysis provides insights into the underlying mechanism, turning $\mathcal{NC}_{1}$ into an indicator for when multi-proxy imprinting is highly benefitial.
@@
Our experiments are limited to foundation models for image classification and do not cover other data modalities such as audio, text, or sensor inputs. While IMPRINT is agnostic to modality, empirical validation of our results outside of vision is needed. Although imprinting alone provides an efficient solution to transfer learning, when compared to purely gradient-based learning, a gap still remains. We do not investigate the benefit of combining it with optimization methods such as gradient-based learning, and the choice of $k$ still requires heuristic or empirical selection rather than direct prediction.
@@
The usage of both weight and activation sparsity as shen2023reducing could change the within- and between-class variability in favor of using a higher number of proxies. Besides only using the penultimate layer embeddings for generating the classifier weights, an interesting area of study could be extracting embeddings from previous layers of the FM for this purpose. Recently, the study by marczak2025revisiting showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable. Another avenue of research is to thoroughly analyze imprinting the weights of other layers as well (masked-imprinting).
Acknowledgments
Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – FIP-12 – Project-ID 528483508, as well as the European Union under the grant project 101079894 (COMFORT - Improving Urologic Cancer Care with Artificial Intelligence Solutions). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. We thank Viet Anh Khoa Tran for initial discussions about the neural collapse phenomenon. We further thank the reviewers and the Action Editor for their detailed and constructive feedback, which substantially improved this work.
@@ Appendix A Appendix
A.1 Additional Results
We provide additional tables and critical difference (CD) diagrams that expand on the results of section˜5.
@@ A.1 Additional Results
[table]A table beside a figure GEN $k$ Avg. acc. % $\pm$ std k-means 5 89.15 $\pm 6.96$ mean 1 86.84 $\pm 7.80$ k-medoids 5 84.87 $\pm 8.76$ k-cov-max 5 82.11 $\pm 10.39$ k-random 5 75.93 $\pm 11.92$ k-fps 5 63.64 $\pm 12.29$ Refer to caption
@@
Figure A.1: Benchmarking GEN mechanism for $k\leq 5$ across FMs and Ts. Best NORM combination for each row is used implicitly. AGG is fixed to max. CD diagram depicts statistical significance of k-means as GEN. See fig.˜5 for $k\leq 20$.
[…]
[table]A table beside a figure NORMinf NORMpre NORMpost Avg. acc. % $\pm$ std L2 none L2 86.84 $\pm 7.80$ none none L2 86.84 $\pm 7.80$ L2 L2 L2 86.79 $\pm 7.83$ L2 none quantile 82.90 $\pm 12.87$ none none quantile 82.90 $\pm 12.87$ L2 L2 quantile 82.83 $\pm 12.86$ L2 L2 none 83.26 $\pm 9.18$ L2 none none 67.66 $\pm 22.12$ none none none 67.66 $\pm 22.12$ Refer to caption
@@
Figure A.2: Benchmarking NORM across FMs and Ts shows crucial effect of L2 normalization. GEN is fixed to mean and AGG to max. CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as NORMinf & NORMpre & NORMpost.
[…]
[table]A table beside a figure NORMinf NORMpre NORMpost Avg. acc. % $\pm$ std L2 none L2 91.04 $\pm 6.26$ none none L2 91.04 $\pm 6.26$ L2 L2 L2 91.06 $\pm 6.21$ L2 L2 quantile 90.51 $\pm 6.40$ L2 L2 none 89.55 $\pm 6.70$ L2 none quantile 79.21 $\pm 15.20$ none none quantile 79.21 $\pm 15.20$ L2 none none 73.53 $\pm 21.05$ none none none 73.53 $\pm 21.05$ Refer to caption
@@
Figure A.3: Benchmarking NORM across FMs and Ts. GEN is fixed to k-means with $k=20$ and AGG to max. CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as NORMinf & NORMpre & NORMpost.”
@@
Figure A.4: k-means with different values for $k$ in n-shot scenarios with focus on $10\leq n\leq 400$. Shaded areas indicate $95\%$ confidence intervals (CIs) over five different seeds. Other variables are fixed to our previously described best method of using L2 for NORM and max as AGG. Note that only data for the meaningful case of $k\leq n$ is shown. For MNIST and FashionMNIST, mean ceases to be the best strategy already at only roughly 50 samples. See fig.˜10 for more values of $n$.
A.2 Datasets
We briefly describe the datasets used in our experiments.
ImageNet (imagenet).
We use the ILSVRC 2012 version (commonly called ImageNet-1K) containing 1 000 classes and 1.2M training images. Since the test set is not publicly available, we use the validation set (50 000 images) as a stand-in. For neural collapse investigations, we construct tasks by randomly grouping $d$ classes into one label, producing “$d$ in 1” tasks as explained in section˜4.
MNIST (mnist).
A benchmark dataset of handwritten digits (0–9), consisting of 60 000 training and 10 000 test grayscale images of size $28\times 28$.
FashionMNIST (fashionmnist).
A drop-in replacement for MNIST with the same format and number of samples, containing grayscale images of fashion items across 10 classes.
CIFAR-10 (cifar10).
A dataset of $32\times 32$ RGB images covering 10 object classes, with 50 000 training and 10 000 test samples.
MNIST-M (mnistm), SVHN (svhn), USPS (usps).
Digit classification datasets each containing digits 0-9 with domain-specific visual characteristics. MNIST-M applies color and texture transformations to MNIST digits, yielding 60 000 training and 10 000 RGB images of size $28\times 28$. SVHN consists of digit crops from house numbers in Google Street View, totaling 73 257 training and 26 032 test images of $32\times 32$ in RGB. USPS contains scanned and normalized handwritten digits in $16\times 16$ grayscale, split into 7 219 training and 2 007 test images.
CombiDigits (Ours).
A synthetic dataset constructed by merging all classes from MNIST, MNIST-M, SVHN, and USPS. Each class label corresponds to a digit (0–9) and fig.˜A.5 shows example images of class “6”. The dataset includes significant visual heterogeneity across sources and thus simulates multi-modal, non-collapsed class distributions.
Figure A.5: Twenty sample images from class “6” of the CombiDigits dataset. Each row corresponds to one of the four source datasets: MNIST, MNIST-M, SVHN, and USPS (top to bottom), illustrating the intra-class heterogeneity across domains.
A.3 Imprinting as Memory
We revisit the classical idea of Bidirectional Associative Memories (BAMs) (kohonen; anderson72; nakano; anderson77) and the associated update rule (sejnowski; dayan) for storing key-value pairs as in
$\mathbf{W}\leftarrow\mathbf{W}+\boldsymbol{\mathbf{v}}\boldsymbol{\mathbf{k}}^{\top},$
where $\mathbf{W}\in\mathbb{R}^{l\times l}$ is a matrix and $\boldsymbol{\mathbf{k}},\boldsymbol{\mathbf{v}}\in\mathbb{R}^{d}$ are key and value vectors, respectively, to be stored therein.
Throughout this work, imprinting can be interpreted as inserting such key-value associations into the classifier weights. More precisely, the generation (GEN) component extends the linear classification head by setting $\boldsymbol{\mathbf{v}}$ as a one-hot vector representing the class, and $\boldsymbol{\mathbf{k}}$ as the corresponding proxy previously computed. While we focus on this simple linear setting, imprinting is not limited to it, as discussed in section˜2.
Notably, this form of direct memory update has seen renewed attention in modern architectures beyond standard query-key-value attention. In particular, the xLSTM model (xlstm) implements this mechanism within its mLSTM memory blocks, where the matrix memory cell is updated by gated key-value associations, closely following this classical covariance rule, indicating a broader resurgence of associative memory principles in contemporary sequence modeling.
A.4 Differences between Foundation Models
@@
While an in-depth comparison of foundation models is beyond the scope of this work, we believe it is important to highlight key observations we have made. In particular, fig.˜11(a) shows significantly lower $\mathcal{NC}_{1}$ scores for vit_b_16 and swin_b on their pre-training ImageNet data compared to the resnet models. We hypothesize that this difference is primarily due to model size and training regimes. The Transformer-based architectures (vit_b_16 and swin_b) have a considerably higher parameter count ($\approx 87$M) than the resnet models (11.7M and 25.6M, respectively). Additionally, vit_b_16 and swin_b were trained for more than three times as many epochs (300 vs. 90) while using a substantially lower learning rate (0.003 and 0.01 vs. 0.1). Notably, the embedding dimensions of these models are comparable, meaning that the observed differences in $\mathcal{NC}_{1}$ scores cannot be attributed to differences in representation dimensionality. Instead, we argue that the combination of larger model size, extended training duration, and lower learning rates likely contributes to greater overfitting, leading to more pronounced collapse. As discussed in section˜5.3, this enables the Transformer-based FMs to handle the ImageNet tasks with remapped labels more effectively and to achieve significantly better performance on the similarly distributed CIFAR-10 dataset.
(a) k-means used as GEN. NORMpost and NORMinf are set to L2, and NORMpre to none. Besides the prominent peaks in accuracy at $k=d$ (as already observed in fig.˜11(b)), a consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, which was not to be seen with NORMpre set to L2 as well. In section˜A.4 we hypothesize that this is due to the distinct embeddings distributions of CNN- and Transformer-based architectures.
(b) k-least-squares (see section˜A.5) used as GEN and all NORM are set to none. Here, the CNN-based FMs exhibit a clear accuracy peak at $k=d$, with a consistent drop for intermediate values $1<k<d$. For the Transformer-based FMs, however, accuracy is typically highest at $k=1$ and only improves slightly around $k=d$ but does not show a pronounced peak. For all FMs, performance declines with $k>d$, indicating that excessive proxy splitting harms generalization. This degradation does not occur in k-means GEN, which remains robust even with large $k$ (see figs.˜11(b) and 6(a)).
@@
Figure A.6: Averaged accuracy of ten random ImageNet label remappings ($d$ in 11) for every $d=1,\dotsc,10$ over number of proxies $k$. Shaded areas indicate $95\%$ CIs over three different seeds. Values for the tasks containing all of MNIST, FashionMNIST, CIFAR-10, resp. CombiDigits at once are shown in dotted lines.
@@
Figure˜6(a), similar to fig.˜11(b), illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs). The key difference in this figure is the use of none for NORMpre instead of L2. This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig.˜11(b), where L2 is used as NORMpre, and does not occur at all in the resnet models. We hypothesize that this difference arises from the distinct embedding distributions of CNN- and Transformer-based architectures (see, e.g., itsdone).
A.5 Learned Weights (with Multiple Proxies) and Comparisons
harun2025good recently studied the initialization of classifier weights for novel categories in a Continual Learning (CL) setting using the least squares algorithm, comparing it to random initialization and class mean imprinting across various loss functions. In contrast, our work centers on imprinting-based approaches, which avoid using gradient-based optimization and cross-class statistics, operating instead on a per-class basis with immediate availability of data. Nonetheless, we include the least-squares method as a supervised oracle baseline – explicitly not an imprinting method – fine-tuned for classification accuracy.
We evaluate the performance of least-squares on the same tasks as defined in section˜4. Additionally, we introduce a multi-proxy extension, k-least-squares, which combines least squares with k-means clustering. As in the case of weight generation GEN in imprinting (where k-means outperforms mean), we find that using multiple proxies per class improves performance in settings with high $\mathcal{NC}_{1}$ as well, indicating that the benefits of multi-prototype representations persist even in this non-imprinting, supervised context.
All experiments use no normalization (none as NORM), based on ablations confirming that additional normalization impairs performance. This matches expectations, since least squares outputs are already calibrated, and normalization distorts them.
Least Squares Weights.
For all data contained in task T, define $\mathbf{H}\in\mathbb{R}^{l\times N}$ as the collection of all feature vectors of the $N$ training samples in T obtained by applying a fixed FM, i.e., $\boldsymbol{\mathbf{h}}_{c,i}\in\mathbb{R}^{l}$ is the feature vector of the $i$-th sample in the $c$-th class. Recall from section˜3.2 that the within-class covariance matrix $\mathbf{\Sigma}_{W}$ is given as $\mathbb{E}_{c,i}[(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\mathbf{h}}_{c})(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\mathbf{h}}_{c})^{\top}]$ and additionally define the total covariance matrix $\mathbf{\Sigma}_{T}$ and the class-means matrix $\mathbf{M}$ as
$\mathbf{M}=[\overline{\boldsymbol{\mathbf{h}}}_{1},\ldots,\overline{\boldsymbol{\mathbf{h}}}_{C}]\in\mathbb{R}^{l\times C},\qquad\mathbf{\Sigma}_{T}=\mathbb{E}_{c,i}[(\boldsymbol{\mathbf{h}}_{c,i}-\boldsymbol{\mathbf{h}}_{G})(\boldsymbol{\mathbf{h}}_{c,i}-\boldsymbol{\mathbf{h}}_{G})^{\top}]\in\mathbb{R}^{l\times l}.$
From these feature statistics, we can obtain the least-squares weights via
$\mathbf{W}_{LS}=\frac{1}{C}\mathbf{M}^{\top}(\mathbf{\Sigma}_{T}+\boldsymbol{\mathbf{h}}_{G}\boldsymbol{\mathbf{h}}_{G}^{\top}+\lambda\mathbf{I})^{-1}.$ (A.1)
Here, $\mathbf{I}$ is the identity matrix and $\lambda$ is the weight decay. We set $\lambda$ to match the value used during the original training of each respective model. Table˜A.1 lists the specific $\lambda$ values used for all models considered in our experiments.
Table A.1: Weight decay values $\lambda$ used for each model.
Model $\lambda$
resnet18, resnet50 (resnet) 0.0001
vit_b_16 (vit) 0.1
swin_b (swin) 0.05
Equation˜A.1 reveals that least-squares shares structural similarities with mean imprinting, but applies a more sophisticated normalization scheme that includes both scaling and rotation. Crucially, it relies on cross-class statistics computed across the entire dataset, and thus is incompatible with imprinting scenarios, which operate by directly constructing class weights from the data of a single class without supervision or access to other classes.
Algorithm 1 k-least-squares
1:Class data $\{\mathbf{H}_{c}\}_{c=1}^{C}$, number of proxies $k$
2:Weights $\{\mathbf{W}_{c}\}_{c=1}^{C}$ with shape $[k,l]$ per class
3:if $k=1$ then
4:  return standard least squares weights $\mathbf{W}_{LS}$ for each class (see eq.˜A.1)
5:end if
6:for each class $c$ do
7:  Cluster $\mathbf{H}_{c}$ into $k$ clusters via $k$-means
8:  Assign each cluster $j$ to proxy class $(c,j)$
9:  Collect proxy samples $\{\mathbf{H}_{(c,j)}\}_{j=1}^{k}$
10:end for
11:Compute least squares weights $\boldsymbol{\mathbf{w}}_{(c,j)}$ jointly for all proxy classes $(c,j)$
12:Assemble $\mathbf{W}_{c}=[\boldsymbol{\mathbf{w}}_{(c,1)},\ldots,\boldsymbol{\mathbf{w}}_{(c,k)}]$ for each original class $c$
13:return $\{\mathbf{W}_{c}\}_{c=1}^{C}$
Combining k-means and least-squares into k-least-squares.
To allow for multiple proxies per class, we propose k-least-squares, a generalization of standard least squares (least-squares) that integrates clustering. Instead of computing a single weight vector per class using all class samples, we first partition each class’s feature set $\mathbf{H}_{c}$ into $k$ clusters using $k$-means. Each cluster is treated as a separate proxy class, effectively expanding the classification task from $C$ to $k\cdot C$ proxy classes. We then solve a single regularized least squares problem over all proxy classes at once, assigning a distinct target vector to each proxy. The resulting weights are grouped by their original class to yield $k$ weight vectors per class. The complete procedure is given in algorithm˜1.
The results in fig.˜6(b) extend previous findings from mean and k-means for the CNN-based FMs, as accuracy increases with increasing proxy count $k$, peaks at $k=d$, and reveals a distinct dip between $k=1$ and $k=d$, confirming the weakness of using only one proxy per class. For Transformer-based FMs, by contrast, accuracy is often highest at $k=1$, improves slightly near $k=d$, but lacks a pronounced peak. Beyond $k>d$, accuracy generally declines for k-least-squares, indicating diminishing returns from excessive proxy splitting. In contrast, k-means GEN maintains stable performance even as $k$ grows, demonstrating robustness to increasing proxy counts.
Comparison with k-means.
We now compare the performance of k-means and k-least-squares as proxy generation methods, noting again that k-least-squares is not an imprinting scheme as it does not operate on a class-by-class basis as the imprinting methods presented in section˜3.1 do.
Upon proper comparison of the best performing k-means and k-least-squares configurations, as depicted in fig.˜A.7, we observe that on our synthetic ImageNet tasks, k-least-squares does not consistently offer a substantial improvement. In fact, it is only on resnet50, in a highly multi-modal setting ($d=10$), that k-least-squares reaches better accuracy. For the Transformer-based models, k-means even performs better in these scenarios.
While the Transformer-based models generally achieve superior performance overall (note the different y-axis scales for CNN- and Transformer-based models in fig.˜A.7), a consistent trend across all models is observed: with increasing multi-modality (increasing $d$), using more than one proxy, whether through k-means or k-least-squares, begins to outperform the single-proxy methods (mean and least-squares). Still, it is striking to see how effective single weights with least-squares can be. However, it is also noteworthy how easily k-means can bridge this gap, particularly considering that least-squares is not an actual imprinting scheme, while k-means is.
Figure A.7: Averaged accuracy of ten random ImageNet label remappings ($d$ in 11) for every $d=1,\dotsc,10$ for mean and least-squares, and optimal k-means and k-least-squares ($k\in\{0,\dotsc,15\}$) weights. Shaded areas indicate $95\%$ CIs over three different seeds. Note the differing y-axis scales for CNN-based (resnet18, resnet50) and Transformer-based (vit_b_16, swin_b) models. The figure illustrates that with increasing multi-modality $d$, multi-proxy methods (k-means, k-least-squares) generally outperform single-proxy methods (mean, least-squares), and that the k-means imprinting scheme is competitive with least-squares approaches.
Furthermore, we evaluate the k-least-squares weight generation method (using none for all NORM) across the 48 tasks proposed in the main part of our paper (see section˜4). least-squares reaches an average accuracy of 94.5494.54%, while 20-least-squares drops to 91.2091.20%. The observed decrease in accuracy can be explained by the dynamics shown in fig.˜6(b). In comparison to the numbers presented in fig.˜2, this shows that our $k$-means imprinting scheme significantly bridges the gap between single-proxy imprinting ($k=1$) and the optimal least squares weights.
A.6 Computational Efficiency
Clustering-based imprinting.
Our k-means implementation in GEN uses sklearn.cluster.KMeans with its default parameters. The cost is $\mathcal{O}(Nklt)$ for $N$ samples, $k$ clusters, feature dimension $l$, and $t$ Lloyd iterations. While $k$-means is not computationally negligible, both assignment and update steps in each iteration parallelize naturally, and convergence typically requires only a few iterations. Similarly, the other steps in our imprinting framework (NORM, AGG) scale linearly with dataset size or proxy count and parallelize efficiently, resulting in no practical scalability bottlenecks even for larger or more complex datasets.
Gradient-based optimization.
The closed-form least-squares costs $\mathcal{O}(Nl^{2}+l^{3}+NlC)$ from the covariance computation and matrix inversion. If solved through stochastic gradient descent, the cost is $\mathcal{O}(NlCt)$ with $t$ epochs.
Empirical runtime.
Table˜A.2 shows that least-squares is faster than k-means in practice, likely because it processes all data in a single closed-form step rather than iterating class-by-class. Nevertheless, it should be clarified again that least-squares represents the analytic, non-iterative optimum derived from gradient minimization and lacks sequential class handling – precisely where imprinting excels (e.g., in continual or edge-learning scenarios).
Table A.2: Average runtime in seconds for different GEN variants across all 48 tasks (see section˜4). To ensure transparency, all timing measurements were obtained on identical hardware (8 vCPUs on Intel Xeon Gold 6438Y+ nodes (2 sockets, 64 physical cores/128 threads)). Cores were not pinned to processes, which may introduce minor variance; however, we observed less than 10% variation across seeds.
GEN variant Runtime (s)
mean 0.0029
5-means 1.2249
20-means 1.7321
least-squares 0.1593
5-least-squares 1.3724
20-least-squares 2.3717